3. Find the slope and intercept of the line x – 2y = 3.
Answer by stanbon(75887)

Find the slope and intercept of the line x – 2y = 3.
-----------------------------------------
Write in slope-intercept form.
2y = x-3
y = (1/2)x - (3/2)
--------------------------
slope = 1/2
intercpt = -3/2
=============================
Cheers,
Stan H.